If ‘S’ is stress and ‘Y’ is young’s modulus of material of a wire, the energy stored in the wire per unit volume is
(a) S2/2Y
(b) 2S2Y
(c) S/2Y
(d) 2Y/S2

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
. (a) Energy stored per unit volume,
E = 1/2 * stress * strain
We know that,
E = 1/2 * stress * stress/Y = 1/2 .S2/Y